# Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection

### v1

@@ 4. Experiments
[…]
Det.(tune)
[…]
Knowledge
[…]
4.1. Metrics and Settings
FactCls Metric
We employ the FactCls metric, denoted by the Micro F1 score, to assess binary factuality classification performance, with a distinct focus on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual. In light of our commitment to pinpointing non-factual examples, we designate non-factual as the positive class and factual as the negative class.
ExpMatch Metric
In FactCHD, the chain of evidence integrates two elements: introductory, expository statements and a core of factual explanations. The former provides phrasing like ’Therefore, there is an incorrect conclusion in this query and response’ as a segue, while the latter delivers detailed insights, bolstering the reasoning behind factuality judgments. Given the paramount importance of aligning factual explanations over expository parts, we employ a segmented matching and weighted averaging metric, ExpMatch, amalgamating these sections for evaluative precision. Specifically, we determine $\text{Score}_{bd}$ using span-based Micro F1 to measure word overlap at the unigram level between the model-generated and reference bodies, reflecting factual core matching. Concurrently, $\text{Score}_{ht}$ is defined via the ROUGE-L score for the head-to-tail part, assessing similarity based on the longest common word subsequence between system-generated context and gold reference. Ultimately, we introduce the ExpMatch metric, a composite of $\text{Score}_{bd}$ and $\text{Score}_{ht}$, serving as the pivotal metric for appraising the coherence and quality of utilized factual chains in explanations:
(1) $\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$
In our experiment, we initially assign $\alpha$ to 0.7. If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation.
4.2. Empirical Examination of Baseline Strategies
This study seeks to probe the efficiency and viability of several prototypical baseline strategies in the detection of fact-conflicting hallucinations. These strategies encompass: (1) Zero-shot Learning, which assesses model performance without prior training in hallucination detection; (2) In-context Learning, employing 4-shot samples as demonstrations to prompt the model and evaluate its competence in managing context-specific hallucinations; (3) Detect-Specific Expert Model, finessing the LoRA (Hu et al., 2022) parameters through fine-tuning, utilizing our specialized training set to harness domain expertise; (4) Knowledge Enhancement, amplifying the model’s capabilities by integrating external knowledge via retrieval techniques or tool-based enhancements. Through these experiments, we aim to traverse both the potentials and limitations of these baseline methods within the scope of hallucination detection.
@@ 4.2. Empirical Examination of Baseline Strategies
We evaluate the performance of various prominent LLMs within the context of our FactCHD benchmark, primarily focusing on models accessible through the OpenAI API, such as text-davinci-003, which corresponds to the InstructGPT model trained using supervised instruction tuning (Ouyang et al., 2022), as well as GPT-3.5-turbo (ChatGPT), which represents one of the most widely recognized models. Alternatively, we explore the adoption of open-source models such as Llama2-chat (Touvron et al., 2023b), Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023), which are fine-tuned variants of the LLaMA (Touvron et al., 2023a).
@@
We conduct the generation of factuality evaluation instances using the Azure OpenAI ChatGPT API [^2] https://openai.com/blog/chatgpt/. During the generation process, we employ a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. Additionally, we set the frequency penalty to zero and utilize a Top-$p$ value of $1.0$ to ensure unrestricted token selection during generation. For evaluation purposes, we set the temperature to 0.2 across all models to reduce output randomness. Specific instructions employed in experiments as detailed in Appendix B.
Figure 4. Overview Truth-Triangulator. Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments. We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
4.2.1. Zero-shot Learning Performance
@@
The empirical results of the evaluation, presented in Table 3, reveal that LLMs with zero-shot learning still struggle with identifying implicit factuality in “Query-Response” contexts. The performance of the open-source 7B LLMs in zero-shot learning is notably poor, highlighting its deficiency in both instruction comprehension and internal knowledge representation. Even the ChatGPT exhibits limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving a mere 52.82% FactCls score in zero-shot, just slightly above chance level. Further investigation into the underwhelming performance reveals a notable drawback: LLMs (especially Llama2-chat) are susceptible to false negatives, resulting in reduced sensitivity when it comes to error detection.
4.2.2. In-context Learning Performance
[…]
4.2.3. Detect-Specific Expert Performance
This study explores the efficacy of tuning LLMs using LoRA to serve as domain-specific experts in identifying fact-conflicting hallucinations. Through training LLMs with our dedicated instructions and datasets, which are intrinsically focused on factuality, we aim to hone their proficiency in this particular domain. Our results highlight a notable enhancement across all open-source 7B models, wherein Llama2-chat emerges as the top-performing model amongst its peers. Post-fine-tuning with fact-conflicting hallucination data, Llama2-chat-7B models attain peak performance on our benchmark, as evidenced by a 74.73% FactCls score and a 53.71% ExpMatch score, respectively. The demonstrated success in training a scaled-down 7B model using our approach underlines the potential of employing 7B LLMs as effective hallucination detectors and paves the way for additional research and trials in this field.
4.2.4. Knowledge Enhancemant
Retrieval Enhancement
@@ 4.2.4. Knowledge Enhancemant
Leveraging retrieval of pertinent knowledge stands as a prevalent strategy to mitigate factual inaccuracies (Lewis et al., 2020). This study empowers LLMs with knowledge facts derived from Wikipedia, deploying TF-IDF for document retrieval and subsequently applying BM25 to pinpoint the top two most pertinent paragraphs from the fetched document. Given the 7B model’s constrained capacity to comprehend lengthy instructional texts—as unveiled in our ICL experiments—we elevate its grasp of expansive knowledge by supplying knowledge for fine-tuning the LoRA parameters. However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest. We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia knowledge encompasses merely a subset, and our retrieval method, being elementary and generic, occasionally yields lower-quality evidence.
Table 4. Out-of-distribution case analysis of our Truth-Triangulator.
@@
User Contexts and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
@@
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict:
NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
@@
Tool Enhancement
Considering the labor-intensive nature of tapping into external knowledge bases, we extend our exploration into exploiting ChatGPT’s potent contextual capabilities for tool-enhanced learning. Specifically inspired by preceding works such as (Chern et al., 2023; Qin et al., 2023), we harness the tool’s execution results to bolster large-scale models in hallucination detection. Employing a search engine as a tool to address queries from the “Query-Response” samples, we prompt ChatGPT to ascertain hallucination presence in the input “Query-Response” grounded on the search engine’s retrieved results, thereby providing concomitant justifications. Experimental outcomes, detailed in Table 3, not only showcase this method’s substantial improvement over the zero-shot approach predicated on ChatGPT but also further underscore the promising potential imbued in tool-enhanced learning.
Table 5. Ablation analysis on input context on FactCls.
@@
4.3. A Strong Baseline: Triangulation for Truth
We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge. However, it is inevitable that the information returned by external knowledge sources may be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model. On the other hand, the detect-specific expert as the Truth Guardian relies on its own knowledge and expertise in the task, tending towards more conservative predictions. To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent sources or perspectives. Specifically, we fine-tune ChatGPT with over 10 examples as a Fact Verdict Manager and collect evidence from different angles or viewpoints to enhance the reliability and accuracy of the obtained truth or conclusion. Table 3 demonstrates that our model, compared to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), exhibits improvements. This finding further underscores the effectiveness of triangulation in mitigating errors and inconsistencies that may arise from relying on a single source or method, thereby facilitating a more comprehensive and robust understanding of the truth.
5. Analysis
5.0.1. Examining the Influences of Model Capacity.
As depicted in Figure 5, our insights reveal a notable enhancement in detecting fact-conflicting hallucinations when migrating from 7B to 13B models, especially within the realms of zero-shot and in-context learning scenarios. Notably, Alpaca-13B surpasses ChatGPT in performance, a phenomenon which can be ascribed to the implementation of consistent prompts throughout all experiments and a meticulous adjustment of prompts to cater to Alpaca’s heightened sensitivity in comparison to ChatGPT  [^3] Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal.. Intriguingly, when models are fine-tuned with training data, the impact of model capacity on performance amplification appears to be minimal. This observation intimates that training LLMs to function as hallucination detectors may effectively alleviate the necessity for augmented model capacities, thereby presenting an efficacious alternative pathway for enhancing model performance.
Figure 5. Impact of model size on hallucination detection.
5.0.2. Elevating Detection Capabilities: Analysis on Accurate and Rich Context
As discussed above, the incremental increase in model parameters for improving hallucination detection has shown diminishing returns, even up to the 13B model. To further investigate this, we orchestrate an ablation study aimed at scrutinizing the impact of implementing accurate facts for knowledge enhancement, as well as integrating query-based “Query-Response” context, to amplify the capabilities in hallucination detection. This investigation seeks to elucidate the nuanced role that precise and rich context plays in refining the model’s detection proficiency.
Enhancement lies in Accurate Fact.
We additionally investigate the straightforward application of the dataset’s intrinsic facts (i.e., the facts utilized during dataset construction) as retrieved evidence within the LoRA fine-tuning process. As illustrated in Table  5, incorporating factual information (w/ fact) culminates in a notable elevation in the FactCls score, substantiating the idea that the modest enhancement derived from retrieval is ascribed to the inferior quality of the garnered evidence. This also accentuates the considerable potential for augmentation by seeking the utmost accurate facts for evaluation.
Complete “Query-Response” Context.
A distinct characteristic of our dataset, which sets it apart from traditional fact-checking datasets, involves incorporating the “Query-Response” context. To delve deeper into its significance, we execute experiments, omitting the “query” during the Alpaca-7B-LoRA fine-tuning process. Table 5 displays the outcomes, wherein the “w/o query” scenario leads to a substantial decline of approximately 40% in the FactCls score, underscoring the pivotal role of a comprehensive “Query-Response” context in extracting invaluable information for informed decision-making.
5.0.3. Case Analysis.
To underscore the broad applicability of Truth-Triangulator, we embarked on supplementary tests, applying it to real-world hallucination data that transcends the boundaries of the FactCHD benchmark. We elucidate both the capabilities and constraints of our model by showcasing the findings from our out-of-distribution case analysis in Table 4. The results from these cases substantiate that our method can make proficient judgments, particularly when discrepancies arise between the detect-specific expert and tool-enhanced ChatGPT. This enhances the reliability of fact-conflicting hallucination detection by providing a practical examination of its application in genuine, unscripted scenarios.
6. Conclusion and Future Work

### v2

@@ 4.5 Quality Filtering via Human Annotator
GPT-3.5-turbo 62.02 37.29 65.66 51.85 32.2 48.11 64.74 50.14 $\uparrow$8.2261.04 $\uparrow$17.9341.96
text-davinci-003 56.52 39.36 55.02 58.22 8.50 48.53 50.34 51.82 $\uparrow$1.952.88 $\uparrow$24.8844.45
Alpaca-7B 35.82 31.01 18.12 40.16 8.86 29.28 6.70 31.52 $\uparrow$5.2428.34 $\uparrow$16.7632.23
Vicuna-7B 41.36 42.51 29.24 58.35 19.36 41.55 13.46 53.60 $\uparrow$6.335.14 $\uparrow$19.1445.98
Llama2-7B-chat 31.00 39.08 39.13 54.38 10.50 41.83 27.96 51.73 $\uparrow$24.4830.25 $\uparrow$13.6143.02
Det.(tune)
Alpaca-7B-LoRA 73.14 49.00 63.34 70.83 69.92 59.88 68.18 63.75 $\uparrow$42.3270.66 $\uparrow$22.7354.96
Vicuna-7B-LoRA 73.52 48.07 64.72 71.74 67.34 62.08 50.36 66.04 $\uparrow$34.4469.58 $\uparrow$9.0054.98
Llama2-7B-chat-LoRA 77.41 47.91 67.70 67.30 62.27 57.03 78.68 65.94 $\uparrow$44.4874.73 $\uparrow$10.6953.71
Knowledge
Alpaca-7B-LoRA (wiki) 73.86 49.44 67.3 69.97 68.24 60.25 67.38 63.00 $\uparrow$0.6671.32 $\uparrow$0.1155.07
Vicuna-7B-LoRA (wiki) 75.14 49.56 65.46 72.71 65.10 63.51 55.42 66.65 $\uparrow$1.28 70.86 $\uparrow$1.3056.28
Llama2-7B-chat-LoRA (wiki) 77.14 46.71 69.61 64.17 66.05 49.73 78.08 64.52 $\uparrow$1.13 75.86 $\uparrow$0.8754.58
GPT-3.5-turbo (tool) 69.71 38.60 69.92 48.43 44.08 47.26 74.21 45.65 $\uparrow$7.5968.63 $\uparrow$0.4142.37
@@
5 Experiments
5.1 Metric Definition
FactCls Metric.
We employ the FactCls, denoted by the Micro F1 score, to evaluate binary factuality classification performance. This metric focuses on the distribution $p(l|Q\&R)$, classifying instances as either factual or non-factual. With a specific emphasis on identifying non-factual examples, we designate non-factual as the positive class and factual as the negative class.
ExpMatch Metric.
In FactCHD, the golden evidence chain features introductory/expository statements (head-tail part) and a factual explanation core (body part). The former contextualizes with phrases such as ‘Therefore, there is an incorrect conclusion in this query and response’, and the latter delivers the in-depth, fact-based reasoning process. For hallucination detectors, prompts guide their outputs akin to the gold evidence chain for quality assessment. Given the paramount importance of aligning factual explanations over expository parts, we introduce ExpMatch, a metric employing segmented matching with weighted averaging. It computes $\text{Score}_{bd}$ via span-based Micro F1 for unigram overlap between generated and reference bodies, and $\text{Score}_{ht}$ via ROUGE-L to assess similarity based on the longest common word subsequence for the head-to-tail part. ExpMatch, combining $\text{Score}_{bd}$ and $\text{Score}_{ht}$, evaluates the explanations generated from detectors as:
$\small\textsc{ExpMatch}=\alpha\times\text{Score}_{bd}+(1-\alpha)\times\text{Score}_{ht}.$ (1)
We initially set $\alpha$ to 0.7, emphasizing the body part [^2] Preliminary evaluations reveal that scores with $\alpha$ at 0.75 and 0.8 align with our established conclusions. If a generated result lacks factual or non-factual information, indicating a prediction failure, we attribute an ExpMatch value of 0 to that particular example during metric calculation..
5.2 Experimental Settings
@@
We evaluate various leading LLMs on FactCHD benchmark, focusing on OpenAI API models, including text-davinci-003 (InstructGPT) and GPT-3.5-turbo (ChatGPT). Additionally, we explore the adoption of open-source models such as Llama2-chat et al (2023b), Alpaca Taori et al. (2023) and Vicuna Chiang et al. (2023), which are fine-tuned variants of the LLaMA Touvron et al. (2023).
Figure 5: Overview Truth-Triangulator. Here we designate the Refer to caption as the “Truth Guardian” based on Llama2-7B-chat-LoRA while Refer to caption as the “Truth Seeker” based on GPT-3.5-turbo (tool) in our experiments. We want the Refer to caption “Fact Verdict Manager” to collect evidence from different viewpoints to enhance the reliability and accuracy of the obtained conclusion.
@@
Using Azure’s OpenAI ChatGPT API [^3] https://openai.com/blog/chatgpt/, we generate samples with a temperature of $1.0$ to control the diversity of generated samples, while limiting the maximum number of tokens to $2048$ to ensure concise responses. We use a frequency penalty of zero and a Top-$p$ of $1.0$ to ensure unrestricted token selection during generation. For evaluations, we standardize the temperature at $0.2$ to minimize randomness. Detailed instructions are provided in Table 11 through Table 12 of §B.
Baseline Strategy Settings
This study investigates the effectiveness of various baseline strategies in detecting fact-conflicting hallucinations. These strategies include: (1) Zero-shot Learning, which evaluates model performance without prior training in hallucination detection; (2) In-context Learning, using 4-shot samples as demonstrations to prompt the model and evaluate its ability to handle hallucination detection; (3) Detect-Specific Expert Model, fine-tuning the LoRA Hu et al. (2022) parameters using our specialized training set to leverage domain expertise; (4) Knowledge Enhancement, enhancing the model’s capabilities by integrating external knowledge through retrieval techniques or tool-based enhancements. These experiments aim to explore both the potential and limitations of these baseline methods in hallucination detection.
5.3 Empirical Experiment Results
5.3.1 Zero-shot Learning Performance
@@
The empirical results presented in Table 4 show that LLMs with zero-shot learning struggle to identify implicit factuality in “Query-Response” contexts. The performance of the open-source 7B LLMs in zero-shot learning is notably poor, indicating deficiencies in both instruction comprehension and internal knowledge representation. Even the ChatGPT shows limited proficiency in distinguishing between factual and non-factual “Query-Response” samples, achieving only a 52.82% FactCls score in zero-shot. Llama2-chat is susceptible to false negatives, resulting in reduced sensitivity in hallucination detection.
5.3.2 In-context Learning Performance
The incorporation of few-shot information significantly improves fact-conflicting hallucination detection in the GPT-3.5-turbo, Alpaca-7B, Vicuna-7B, and Llama2-7B-chat models. Compared to models operating without this additional context, these enhancements result in an average increase of approximately $\uparrow$6in the FactCls and $\uparrow$18in ExpMatch scores. However, the impact of integrating few-shot information on text-davinci-003 is relatively modest, indicating its limited proficiency in managing few-shot learning. Furthermore, the improvement from incorporating few-shot information into Llama2-chat is significant, possibly because of the model’s superior contextual learning abilities compared to Alpaca and Vicuna, leading to a better understanding of few-shot demonstrations.
5.3.3 Detect-Specific Expert Performance
We investigate the effectiveness of tuning LLMs with LoRA for domain-specific expertise in detecting fact-conflicting hallucinations. By training these models on our trainset, we enhance their detection task proficiency. Our findings show that all tested open-source 7B models improve, with Llama2-chat outperforming its counterparts. After fine-tuning on hallucination data, Llama2-chat-7B achieves a FactCls score of 74.73% and an ExpMatch score of 53.71% on our benchmark. This success with a 7B model underscores the viability of using such models as hallucination detectors and encourages further exploration in this area.
5.3.4 Knowledge Enhancemant
Retrieval Enhancement.
@@ 5.3.4 Knowledge Enhancemant
This study enhances LLM-based detection with Wikipedia-sourced facts, using BM25 for initial document retrieval to select the top five most relevant paragraphs. Our experiments show that given the 7B model’s limited ability to process long texts, providing knowledge during fine-tuning modestly improves performance. However, compared to the detect-specific expert without knowledge augmentation, the improvement achieved by combining Wikipedia retrieval with fine-tuning the LoRA parameters is relatively modest. We attribute these observations to two primary factors: our dataset spans multiple domains while Wikipedia encompasses merely a subset, and our retrieval method being elementary, occasionally yields lower-quality evidence.
@@
Query and Returned Evidence Detection Opinion from Both Sides Prediction with Triangulation
@@
Truth-Triangulator: As the Fact Verdict Manager, after conducting a thorough analysis of both perspectives and evidence, I have reached the following verdict: NON-FACTUAL. The answer that Queen Urraca of Castile was the mother of Afonso II, the third king of Portugal, is incorrect. The evidence provided shows that Afonso II was the son of King Sancho I and Queen Dulcia, daughter of Ramón Berenguer IV of Barcelona. Therefore, the answer contains false assumptions and incorrect conclusions. The Truth Seeker’s response is supported by the provided evidence, while the Proprietary Model’s response is not accurate as it did not consider the external evidence.
@@
Table 5: Case analysis of out-of-distribution examples from ChatGPT using Truth-Triangulator.
Tool Enhancement.
Considering the labor-intensive nature of tapping into external knowledge bases, we investigate leveraging ChatGPT’s advanced contextual understanding for tool-enhanced hallucination detection, drawing from prior research such as Chern et al. (2023); Qin et al. (2023). By using a search engine to inform ChatGPT’s analysis of “Query-Response” samples, we enable the model to identify hallucinations with supporting evidence from search results. Results in Table 4 demonstrate significant enhancements over ChatGPT’s few-shot performance, highlighting the efficacy of tool learning.
5.4 Exploring Triangulation for Truth
We categorize tool-enhanced ChatGPT as the Truth Seeker, which aims to make informed judgments by seeking external knowledge. However, the information returned by external knowledge sources may inevitably be incomplete, erroneous, or redundant, thus potentially misleading the large-scale model. On the other hand, the detect-specific expert as the Truth Guardian relies on its knowledge and expertise in the task, tending towards more conservative predictions. To address these challenges, we propose the Truth-Triangulator framework inspired by the “Triangulation for Truth” theory, involving verifying and confirming information by cross-referencing multiple independent perspectives. Figure 5 illustrates our approach of fine-tuning ChatGPT as a Fact Verdict Manager, leveraging the output and knowledge from Truth Seeker and Truth Guardian to boost conclusion reliability. Table 4 reveals our model’s superior performance to Lamma2-7b-chat-LoRA and GPT-3.5-turbo (tool), highlighting the benefits of triangulation in reducing single-source error and enhancing truth verification.
5.5 Experimental Analysis
Examining the Influences of Model Capacity.
Figure 6 illustrates that transitioning from 7B to 13B models notably improves the detection of fact-conflicting hallucinations, particularly in zero-shot and in-context learning scenarios. Alpaca-13B outperforms ChatGPT, which can beascribed to the consistently adopted command prompt may be more friendly to Alpaca-13B  [^4] We keep consistent prompts for each LLM throughout all experiments. Given that LLMs are acknowledged for their pronounced sensitivity to prompts, we do not assert that our prompts are universally optimal.. Interestingly, when models are fine-tuned with training data, the impact of model capacity on performance improvement appears minimal. This implies that further training larger LLMs as hallucination detectors may yield limited benefits, and it is necessary to explore alternatives with higher upper limits for enhancing detection performance.
Figure 6: Analysis of model capacity impact.
Enhancement lies in Accurate Evidence.
We also explore the direct use of intrinsic facts from the dataset as golden evidence in the LoRA fine-tuning process. As shown in Table  6, integrating factual information (w/ gold evidence) leads to a significant improvement in the FactCls score, indicating that the modest improvement from retrieval may stem from the lower quality of the obtained evidence. This underscores the potential for substantial improvement by seeking the most accurate facts for evaluation.
Complete “Query-Response” Context.
Our dataset differs from traditional fact-checking datasets in its inclusion of the “Query-Response” context. To examine its impact, we conduct experiments by excluding the “query” during the Alpaca-7B-LoRA fine-tuning process. The results in Table 6 reveal that the “w/o query” scenario leads to a significant 40% decrease in the FactCls score, highlighting the essential role of a comprehensive “Query-Response” context in extracting valuable information for informed decision-making.
@@ 5.5 Experimental Analysis
Table 6: Ablation analysis on input context.
Real-World Application Case Study.
We extended the testing of Truth-Triangulator to real-world instances of ChatGPT-generated hallucinations beyond the FactCHD benchmark to demonstrate its wide-ranging utility. The out-of-distribution case analysis, detailed in Table 5, delineates our model’s strengths and limitations. These real-world tests confirm Truth-Triangulator’s adeptness in making accurate assessments, particularly where expert and augmented ChatGPT evaluations diverge, thus bolstering the credibility of detecting fact-conflicting hallucinations in authentic, uncontrolled settings.
6 Conclusion and Future Work